OpenStudy (narissa):

Simplify (-12x^5y^2+6x^3y+2xy)/2xy

OpenStudy (narissa):

@princeharryyy

OpenStudy (narissa):

@airjordan200

OpenStudy (airjordan200):

um i don't think / this is supposed to be thare

OpenStudy (airjordan200):

if not then it is 4x2y2(1 + 3x2 + -6x4y)

OpenStudy (airjordan200):

\[4 x^{2} y ^{2}(1 + 3x2 + -6x4y) \]

OpenStudy (narissa):

What isnt supposed to be there?

OpenStudy (airjordan200):

this -----> /

OpenStudy (airjordan200):

ik it means dividing but i don't believe we need to divide for this

OpenStudy (narissa):

(-12x^5y^2+6x^3y+2xy)÷ 2xy the answer is -6x^4y+3x^2+1 but im not sure how to get that answer. This is only a practice test so i know all the answer i just need a step by step to solve

OpenStudy (princeharryyy):

you just divide like you do simple division nothing more nothing less.

OpenStudy (airjordan200):

Equation at the end of step 1 : (((0-((12•(x5))•(y2)))+((2•3x3)•y))+2xy) (————————————————————————————————————————•x 2 Step 2 : Equation at the end of step 2 : (((0-((22•3x5)•y2))+(2•3x3y))+2xy) (——————————————————————————————————•x)•y 2 Step 3 : -12x5y2 + 6x3y + 2xy Simplify ———————————————————— 2 Step 4 : Pulling out like terms : 4.1 Pull out like factors : -12x5y2 + 6x3y + 2xy = -2xy • (6x4y - 3x2 - 1) Trying to factor a multi variable polynomial : 4.2 Factoring 6x4y - 3x2 - 1 Try to factor this multi-variable trinomial using trial and error Factorization fails Equation at the end of step 4 : (-xy • (6x4y - 3x2 - 1) • x) • y Step 5 : Multiplying exponential expressions : 5.1 x1 multiplied by x1 = x(1 + 1) = x2 Equation at the end of step 5 : -x2y • (6x4y - 3x2 - 1) • y Step 6 : Multiplying exponential expressions : 6.1 y1 multiplied by y1 = y(1 + 1) = y2 Final result : (6x4y - 3x2 - 1)

OpenStudy (narissa):

@jhonyy9